Find the angle theta such that 0greaterthanorequal theta<2pi and cos(2theta)=2-2sin^2 theta
Answer by lwsshak3(11628)

Find the angle theta such that 0greaterthanorequal theta<2pi and cos(2theta)=2-2sin^2 theta
***
cos(2x)=2-2sin^2(x)
1-2sin^2x=2-2sin^2(x)
1≠2 (no solution)